If k is the constant in the expression x^2 – x + k and the expression has a
value of 0 when 3 is substituted for x, what is the value of the expression when 2 is substituted for x?
(A) -4 (B) -2 (C) 6 (D) 8 (E) Cannot be found.

Answer by stanbon(75887)

If k is the constant in the expression x^2 – x + k and the expression has a
value of 0 when 3 is substituted for x, what is the value of the expression when 2 is substituted for x?
----------------------------------------
Find K
3^2 - 3 + k = 0
9-3 + k = 0
6+k = 0
k = -6
--------
Let x = 2
x^2 - x -6 = ?
2^2-2-6 = ?
? = -4
====================
Ans: -4
====================
Cheers,
Stan H.

(A) -4 (B) -2 (C) 6 (D) 8 (E) Cannot be found.